# Deep Reinforcement Learning for Robotic Bipedal Locomotion: A Brief Survey

### v5

@@ Appendix A Deep reinforcement learning algorithms
The advancement and development of RL is crucial for bipedal locomotion. Specifically, advancements in deep learning provide deep NNs serving as function approximators to empower RL with the capability to handle tasks characterised by high-dimensional and continuous spaces, by efficiently discovering condensed, low-dimensional representations of complex data. In comparison to other robots of different morphologies, such as wheeled robots, bipedal robots feature much higher DoFs and continuously interact with environments, which results in higher requirements for the DRL algorithms. Especially in the legged locomotion field, policy gradient-based algorithms are prevalent in the field of bipedal locomotion.
@@
Designing an effective NN architecture is essential for tackling complex bipedal locomotion tasks. Multi-layer perceptrons (MLP), a fundamental NN, excel in straightforward regression tasks with lower computational resource demands. A comprehensive comparison between MLP and the memory-based NN, LSTM reveals that MLPs have an advantage in convergence speed for tasks [80]. However, LSTM, as a variant of Recurrent Neural Networks (RNN), is adept at processing data associated with time, effectively relating different states across time, and modelling key physical properties vital for periodical gaits [32] and successful sim-to-real transfer in bipedal locomotion. Additionally, Convolutional Neural Networks (CNN) specialise in spatial data processing, particularly for image-related tasks, making them highly suitable for environments where visual perception is crucial. This diverse range of NN architectures highlights the importance of selecting the appropriate model based on the specific requirements of the bipedal locomotion tasks.
@@
Considering DRL algorithms, recent bipedal locomotion studies focus on model-free RL algorithms. Unlike model-based RL, which learns a model of the environment but may inherit biases from simulations that do not accurately reflect real-world conditions, model-free RL directly trains policies through environmental interaction without relying on an explicit environmental model. Although model-free RL requires more computational samples and resources, it can train a more robust policy allowing the robots to travel around challenging environments.
@@
Many sophisticated model-free RL algorithms exist, which can be broadly classified into two categories: policy-based (or policy optimisation) and value-based approaches. Value-based methods e.g. Q-learning, State-Action-Reward-State-Action (SARSA), Deep Q-learning (DQN) [72] only excel in discrete action space and often struggle with high dimensional action space. Q-learning is an off-policy algorithm that directly learns the optimal Q-values, allowing it to derive the best possible actions irrespective of the current policy. SARSA, an on-policy variant, updates its Q-values based on the actual actions taken, making it robust in environments where the policy evolves during learning. DQN extends Q-learning by using deep NNs to approximate Q-values, enabling the algorithm to tackle complex state spaces, though it still faces challenges with high-dimensional action spaces due to the difficulty of accurate value estimation. In contrast, policy-based methods, such as policy gradient, can handle complex tasks but are generally less sample-efficient compared to value-based methods.
@@
More advanced algorithms combine both policy-based methods and value-based methods. Actor-critic (AC) refers to a main idea simultaneously learning both a policy (actor) and a value function (critic), where it owns both advantages of both algorithms [124, 125]. Popular algorithms e.g. Trust region policy optimisation (TRPO) [126] and PPO based on policy-based methods, borrow ideas from AC. Moreover, there are other novel algorithms based on the AC framework, Deep Deterministic Policy Gradient (DDPG) [66], Twin Delayed Deep Deterministic Policy Gradients (TD3) [127], A2C (Advantage Actor-Critic), and A3C (Asynchronous Advantage Actor-Critic) [128], SAC (Soft Actor-Critic) [129]. Each algorithm has its strengths considering different tasks in the bipedal locomotion scenario. There are several key factors to value these algorithms such as: sample efficiency, robustness and generalisation, and implementation challenges. A comparative analysis work [65] illustrates that SAC-based algorithms excel in stability and achieve the highest scores, while their training efficiency significantly trails behind that of PPO that obtain relatively high score.
@@
In [64], PPO demonstrates the robustness and computational economy in complex scenarios, such as bipedal locomotion, utilising fewer resources than TRPO. In terms of training time, PPO is much faster than SAC, and DDPG algorithms [65]. Besides, many works [32, 58, 47] have demonstrated its robustness and ease of implementation and combined with the flexibility to integrate with various NN architectures have made PPO the most popular choice in this field. Various work has demonstrated that PPO can conduct the exploration of walking [32], jumping [48], stair climbing [49], and stepping stones [33], which demonstrates its efficiency, robustness and generalisation.
@@
Additionally, the DDPG algorithm integrates the Actor-Critic framework with DQN to facilitate off-policy training, further optimising sampling efficiency. In some explicit scenarios such as jumping, DDPG shows higher reward and better learning performance than PPO [34, 130]. TD3 is developed based on DDPG, and improve over the performance of the DDPG and SAC [129].
@@
SAC improves exploration through its stochastic policy and entropy-regularised objective, which encourages the agent to maintain randomness in its actions, balancing exploration and exploitation more effectively than DDPG and TD3. Unlike PPO, which is an on-policy algorithm, SAC’s off-policy nature allows it to leverage a replay buffer, reusing past experiences to train without requiring constant interactions with the environment. This, combined with entropy maximisation, enables SAC to achieve faster convergence in complex environments where exploration is essential. SAC is also known for its stability and strong performance across a wide range of tasks [129]. While A2C offers improved efficiency and stability compared to A3C, the asynchronous update mechanism for A3C provides better capabilities for exploration and accelerating learning. Although these algorithms show their advancements, they are more challenging to apply due to algorithms’ complexity compared to PPO.
@@ Appendix B Training Simulation Environment
The development of DRL algorithms and sim-to-real techniques has shown the requirement for high-quality simulators. Creating a reliable simulation environment and conducting RL training is difficult. Through the literature, several simulators are available, including Isaac Gym [67], robocup3D [76], Openai Gym [77], MuJoCo [78], Orbit [131], Brax [132], Isaac Lab [133].
@@
For physics-based simulation, MuJoCo developed by DeepMind, and Gazebo are widely used platforms that support a range of robotics research tasks. NVIDIA’s Isaac Gym, although now deprecated, played an important role as a high-performance GPU-based simulator for training agents in complex environments. Its successors, such as Isaac Lab and Orbit, continue to evolve as modern RL and robotics frameworks.
One of the most crucial things is the parallelisation strategy and GPU simulations. For instance, Isaac Gym was developed to maximise the throughput of physics-based machine learning algorithms with particular emphasis on simulations that require large numbers of environment instances executing in parallel. Running the physics simulation on GPU can result in significant speedups, especially for large scenes with thousands of individual actors.

### v6

@@ Appendix A Deep reinforcement learning algorithms
The advancement and development of RL are crucial for bipedal locomotion. Specifically, advances in deep learning provide deep NNs that serve as function approximators, enabling RL to handle tasks characterised by high-dimensional and continuous spaces by efficiently discovering condensed, low-dimensional representations of complex data. In comparison with other robots of different morphologies, such as wheeled robots, bipedal robots possess far higher DoFs and continuously interact with their environments, which results in greater demands on DRL algorithms. In particular, within legged locomotion, policy-gradient-based algorithms are prevalent in bipedal locomotion research.
@@
Designing an effective NN architecture is essential for tackling complex bipedal locomotion tasks. Multi-Layer Perceptrons (MLPs), a fundamental NN structure, excel in straightforward regression tasks with lower computational resource requirements. A comprehensive comparison between MLPs and the memory-based NN LSTM reveals that MLPs have an advantage in convergence speed for many tasks [2023_Rohan_sim2real_currentdecompose]. However, LSTMs, as variants of Recurrent Neural Networks (RNNs), are adept at processing time-associated data, effectively relating different states across time and modelling key physical properties vital for periodic gaits [2021_siekmann_sim2real_nonreference_perodicreward_DRL_e2e_LSTM_PPO_cassie] and successful sim-to-real transfer in bipedal locomotion. Additionally, Convolutional Neural Networks (CNNs) specialise in spatial data processing, particularly for image-related tasks, making them highly suitable for environments where visual perception is crucial. This diversity of NN architectures highlights the importance of selecting an appropriate model based on the specific requirements of bipedal locomotion tasks.
@@
Considering DRL algorithms, recent bipedal locomotion studies have focused on model-free RL algorithms. Unlike model-based RL, which learns a model of the environment but may inherit biases from simulations that do not accurately reflect real-world conditions, model-free RL directly trains policies through environmental interaction without relying on an explicit environmental model. Although model-free RL requires more computational samples and resources, it can train more robust policies that allow robots to traverse challenging environments.
@@
Many sophisticated model-free RL algorithms exist, which can be broadly classified into two categories: policy-based (or policy optimisation) and value-based approaches. Value-based methods, e.g. Q-learning, State–Action–Reward–State–Action (SARSA), and Deep Q-learning (DQN) [2021_DeepQ_DNQ_stepping_animationlonly_steppingstones], excel only in discrete action spaces and often struggle with high-dimensional action spaces. Q-learning is an off-policy algorithm that directly learns the optimal Q-values, allowing it to derive the best possible actions irrespective of the current policy. SARSA, an on-policy variant, updates its Q-values based on the actual actions taken, making it robust in environments where the policy evolves during learning. DQN extends Q-learning by using deep NNs to approximate Q-values, enabling the algorithm to tackle complex state spaces, though it still faces challenges with high-dimensional action spaces due to difficulties in accurate value estimation. In contrast, policy-based methods, such as policy-gradient techniques, can handle complex tasks but are generally less sample-efficient than value-based methods.
@@
More advanced algorithms combine both policy-based and value-based methods. The Actor–Critic (AC) framework simultaneously learns both a policy (actor) and a value function (critic), combining the advantages of both approaches [2015_continous_control_deep_reinforcement_learning, 2016_guided_learning_of_Control_Graphs_physics-based_characters]. Popular algorithms such as Trust Region Policy Optimisation (TRPO) [2015_TRPO_foundamentalpaper] and PPO, based on policy-based methods, borrow ideas from AC. Moreover, other novel algorithms based on the AC framework include Deep Deterministic Policy Gradient (DDPG) [2023_DDPG_e2e_openaigym_mujoco], Twin Delayed Deep Deterministic Policy Gradients (TD3) [2019_stephen_TD3_instruction], A2C (Advantage Actor–Critic), A3C (Asynchronous Advantage Actor–Critic) [2022_Jie_A3C_sim2real], and SAC (Soft Actor–Critic) [2022_Chen_transition_quadtobiped_TD3_SAC_Residual]. Each algorithm has its strengths for different tasks in bipedal locomotion scenarios. Several key factors determine their performance, such as sample efficiency, robustness and generalisation, and implementation complexity. A comparative analysis [2023_Omur_comparison_RLalgorithm_bipedallocomotion] illustrates that SAC-based algorithms excel in stability and achieve the highest scores, while their training efficiency significantly trails behind that of PPO, which attains relatively high scores.
@@
In [2017_Schulman_PPO_instructure], PPO demonstrates robustness and computational efficiency in complex scenarios such as bipedal locomotion, utilising fewer resources than TRPO. In terms of training time, PPO is much faster than SAC and DDPG [2023_Omur_comparison_RLalgorithm_bipedallocomotion]. Moreover, many studies [2021_siekmann_sim2real_nonreference_perodicreward_DRL_e2e_LSTM_PPO_cassie, 2021_Kevin_gaitlibrary_heirarchy_reducedorder_gaitlibrary_sim2real, 2020_Siekmann_drl_e2e_s2r_cassie] have demonstrated its robustness and ease of implementation. Combined with its flexibility to integrate with various NN architectures, this has made PPO the most popular choice in the field. Numerous studies have shown that PPO can enable the exploration of walking [2021_siekmann_sim2real_nonreference_perodicreward_DRL_e2e_LSTM_PPO_cassie], jumping [2023_UCB_Cassie_RL_Jumping], stair climbing [2021_siekmann_blind_DRL_stair], and stepping-stone traversal [2022_OSUDRL_steppingstone_referencefree_predictionfeasiblefootsteps_camera_benchmark], demonstrating its efficiency, robustness, and generalisation.
@@
Additionally, the DDPG algorithm integrates the Actor–Critic framework with DQN to facilitate off-policy training, further optimising sample efficiency. In certain scenarios, such as jumping, DDPG shows higher rewards and better learning performance than PPO [2023_Chongben_jumping_DDPG_multi-task_network_simulationonly, 2022_Chongben_pallel_DDPG_e23_silumation_only]. TD3, developed from DDPG, improves upon the performance of both DDPG and SAC [2022_Chen_transition_quadtobiped_TD3_SAC_Residual].
@@
SAC improves exploration through its stochastic policy and entropy-regularised objective, which encourages the agent to maintain randomness in its actions, balancing exploration and exploitation more effectively than DDPG and TD3. Unlike PPO, which is an on-policy algorithm, SAC’s off-policy nature allows it to leverage a replay buffer, reusing past experiences for training without requiring constant interaction with the environment. This, combined with entropy maximisation, enables SAC to achieve faster convergence in complex environments where exploration is essential. SAC is also known for its stability and strong performance across a wide range of tasks [2022_Chen_transition_quadtobiped_TD3_SAC_Residual]. While A2C offers improved efficiency and stability compared with A3C, the asynchronous update mechanism of A3C provides better exploration capability and accelerates learning. Although these algorithms demonstrate clear advantages, they are more challenging to apply owing to their complexity compared with PPO.
@@ Appendix B Training Simulation Environment
The development of DRL algorithms and sim-to-real techniques highlights the requirement for high-quality simulators. Creating a reliable simulation environment and conducting RL training is challenging. The literature shows that several simulators are available, including Isaac Gym [2021_Sim_Isaac], RoboCup3D [2003_Sim_Robocup], OpenAI Gym [2016_Sim_OpenAI], MuJoCo [2012_Sim_Mujoco], Orbit [2023_Orbit_simulator], Brax [2021_Brax], and Isaac Lab [isaaclab2023].
@@
For physics-based simulation, MuJoCo, developed by DeepMind, and Gazebo are widely used platforms that support a range of robotics research tasks. NVIDIA’s Isaac Gym, although now deprecated, played an important role as a high-performance GPU-based simulator for training agents in complex environments. Its successors, such as Isaac Lab and Orbit, continue to evolve as modern RL and robotics frameworks.
One of the most crucial aspects is the parallelisation strategy and GPU simulation. For instance, Isaac Gym was developed to maximise the throughput of physics-based machine learning algorithms, with particular emphasis on simulations requiring large numbers of environment instances executing in parallel. Running the physics simulation on a GPU can result in significant speed-ups, especially for large scenes with thousands of individual actors.